# Collisionless shocks having relativistic velocities in relativistically hot plasmas

### v1

@@ Abstract
Collisionless shocks play an important role in the dissipation of kinetic energy in various astrophysical objects such as the Sun, supernova remnants, galaxy clusters, black holes, and neutron stars. The dissipation in a collsionless shock converts the upstream kinetic energy to the downstream thermal energy of particles, electromagnetic fields, and nonthermal high energy particles (e.g. Quest, 1988; Shimada & Hoshino, 2000; Spitkovsky, 2008). The some fraction of the thermal and nonthermal particle energy is eventually converted to photons, neutrino, and cosmic ray energies.
@@
The dissipation in a collsionless shock depends on the upstream parameters, such as the sound Mach number, Alfvén Mach number (Gargaté & Spitkovsky, 2012), upstream velocity over the speed of light (Jikei & Amano, 2024), ionization fraction (Ohira, 2013), angle of the magnetic field (Caprioli & Spitkovsky, 2014), and particle composition (Hoshino et al., 1992). Despite the fact that there are many studies about collisionless shocks, collsionless shocks in relativistically hot plasmas ($T\gg mc^{2}$, where $T$, $m$, and $c$ are the upstream temperature, the particle mass, and the speed of light) have been rarely investigated. However, such collisionless shocks are thought to exist in the high energy astrophysical objects, such as the prompt emission region of gamma-ray bursts (Piran, 2004; Kumar & Zhang, 2015), downstream of relativistic shocks propagating in a inhomogeneous medium (Inoue et al., 2011; Tomita et al., 2022), and lobes of radio galaxies (Matthews et al., 2019).
@@
To understand the dissipation in a colissionless shock in a relativistically hot plasma, as a first step, we investigated collisionless shocks propagating with the nonrelativistic and mildly relativistic velocity in a relativistically hot unmagnetized electron-positron plasma using two-dimensional particle-in-cell (PIC) simulations (Kamiido & Ohira, 2024). Our simulation showed that shocks are mediated by the Weibel instability (Weibel, 1959), the strength of the Weibel magnetic fields becomes smaller as the shock velocity becomes smaller, and particle acceleration by the shock is not observed in the simulation time. In Kamiido & Ohira (2024), we discussed the reason for this shock velocity dependence in terms of the dissipation efficiency at the shock front, but did not discuss it in terms of the plasma instability. In addition, the temperature dependence and the case with a relativistic shock velocity were not investigated. In this study, we investigate collisionless shocks propagating with a relativistic shock velocity in relativistically hot unmagnetized electron-positron plasmas. The simulation setup and results are presented in Sections 2 and 3, respectively. To understand the simulation results, we calculate the pressure anisotropy in the shock transition region in Section 4. The discussion and summary are given in Section 5. In the appendices A, B, and C, we derive some physical quantities for the relativistic plasma.
@@ 2 Simulation setup
$\langle\gamma\rangle$ is the mean Lorentz factor of particles in the simulation frame and given by (Equation A2)
@@ 3.2 Magnetic field distribution
Figure 3 shows the distribution of the $z$-component of the magnetic field at $t\omega_{\mathrm{p}}=2000$, which is normalized by the upstream energy density $n_{\mathrm{up}}\langle\gamma\rangle mc^{2}$. The maximum value and typical wavelength of the magnetic field are $B_{z}/\sqrt{8\pi n_{\mathrm{up}}\langle\gamma\rangle mc^{2}}\sim 0.3$ and $\lambda\sim 8\,c/\omega_{\mathrm{p}}$ in common for all simulations, which are almost identical to the results in the case of cold upstream (Spitkovsky, 2008). In addition, the decay rate of the downstream magnetic field is also common to all simulations. Since the Weibel-like magnetic field structure is observed in all simulations, our simulations show that the relativistic unmagnetized collisionless shock is mediated by the Weibel instability, even though the upstream temperature is relativistically hot.
@@ 3.3 Energy spectrum
Figure 4 shows the energy spectrum in the downstream region of $800500<x/(c/\omega_{\mathrm{p}})<$ 800 at $t\omega_{\mathrm{p}}=2000$. As one can see, the energy spectra have the power-law tails with the spectral index of about 2.4 for all simulations, that is, our simulations show that, as in the early studies for the relativistic Weibel mediated shocks in cold plasmas (Spitkovsky, 2008; Sironi et al., 2013), particles are accelerated by the relativistic unmagnetized collisionless shock in the relativistically hot electron-positron plasma.
@@ 4 Pressure anisotropy
Figure 6: Distribution of counter streaming plasmas in the four velocity space. The four velocities $\mathbf{u}$ are normalized by $\langle\gamma\rangle$. The white lines show the cone of the relativistic beaming effect, $u_{y}/u_{x}=\pm 1/\Gamma$, where $\Gamma=20$ is the bulk Lorentz factor corresponding to the counter streaming velocity.
@@
The plasma in the shock transition region consists mainly of two components: the upstream plasma moving from upstream to downstream and the plasma leaking from downstream to upstream. As a simple system, we consider counter streaming plasmas with the bulk velocities in the $x$ direction, ±$\pm\beta c$ and temperature, $T$. The momentum distribution of the counter streaming plasmas is assumed to be the shifted Maxwell-Jüttner distribution function (Equation B2). Then, the pressure anisotropy of the counter streaming plasmas is given by (Equation B3),
@@
Figure 6 shows the momentum distribution in the $u_{x}-u_{y}$ plane of the counter streaming plasmas with $\Gamma=20$ and $T/mc^{2}=0.01,5,500$, where $u_{x}$ and $u_{y}$ are the four velocity in the $x$ and $y$ directions, and normalized by $\langle\gamma\rangle$. $\langle\gamma\rangle$ is the mean Lorentz factor of particles in this system (Equation A2). As shown in the middle and bottom panels in Figure 6, the normalized momentum distribution does not depend on the temperature for the relativistic temperature. Since almost all particles have velocities very close to the speed of light in the each plasma rest frame for the relativistic temperature, in the counter streaming frame, their angular distribution is concentrated in the flow direction in the angle of $\Gamma^{-1}$. This is the so-called relativistic beaming effect. Then, the angular distribution of the counter streaming plasmas is almost independent of the temperature. This is the reason why the pressure anisotropy does not depend on the relativistic temperature.
@@ Appendix A Plasma with a relativistic temperature
where $e$ and $m$ are the charge and mass of each particle.

### v2

@@ Abstract
Collisionless shocks play an important role in the dissipation of kinetic energy in various astrophysical objects such as the Sun, supernova remnants, galaxy clusters, black holes, and neutron stars. The dissipation in a collisionless shock converts the upstream kinetic energy to the downstream thermal energy of particles, electromagnetic fields, and nonthermal high energy particles (e.g. Quest, 1988; Shimada & Hoshino, 2000; Spitkovsky, 2008b). The some fraction of the thermal and nonthermal particle energy is eventually converted to photons, neutrino, and cosmic ray energies.
The Weibel mediated collisionless shock has been actively studied by particle-in-cell (PIC) simulations because the simulation time is shorter than that for other types of collisionless shocks. In the Weibel mediated shock, the upstream plasma is thermalized by magnetic field turbulence generated by the Weibel instability (Weibel, 1959). This instability arises from anisotropic velocity distributions, leading to the generation of filamentary current structures and magnetic fields. PIC simulations showed that Weibel-mediated relativistic shocks are formed in unmagnetized plasmas (Kato, 2007; Spitkovsky, 2008b), particles are accelerated by the shock (Spitkovsky, 2008b; Martins et al., 2009), and electrons are efficiently heated to near ion temperature (Spitkovsky, 2008a). The electron heating in the Weibel mediated shock is theoretically well explained (Vanthieghem et al., 2022). Recent work by Parsons et al. (2024) further examined the injection mechanism to the particle acceleration before particles enter the downstream region. Sironi et al. (2013) established a critical threshold for Weibel-mediated shock formation, showing that such shocks emerge when the upstream magnetic energy fraction satisfies $B_{\mathrm{up}}^{2}/4\pi\Gamma_{\mathrm{up}}n_{\mathrm{up}}mc^{2}\lesssim 10^{-3}$ where $B_{\mathrm{up}}$, $\Gamma_{\mathrm{up}}$, $n_{\mathrm{up}}$, $m$, and $c$ are the upstream magnetic field, the upstream Lorentz factor, the upstream number density at the downstream rest frame, the rest mass of a particle, and the light speed, respectively. In addition, their study showed that particle acceleration in the Weibel mediated shock is insufficient compared to the prediction in the Bohm limit diffusion. Moreover, Lemoine et al. (2019) and Pelletier et al. (2019) developed a theoretical model describing how the interaction between self-generated electromagnetic fields and incoming plasma streams shapes the shock profile. These results about the Weibel mediated shock are well summarized in Vanthieghem et al. (2020). Recently, Grošelj et al. (2024) performed a long-term simulation and showed that the typical length scale of the magnetic field turbulence increases with time, so that the decay of the turbulent magnetic field becomes weaker in the downstream region.
@@
Despite the fact that there are many studies about collisionless shocks, ollisionless shocks in relativistically hot plasmas ($T\gg mc^{2}$, where $T$ is the upstream temperature and we define the Boltzman constant $k_{\mathrm{B}}=1$) have been rarely investigated. However, such collisionless shocks are thought to exist in the high energy astrophysical objects, such as the prompt emission region of gamma-ray bursts (Piran, 2004; Kumar & Zhang, 2015), downstream of relativistic shocks propagating in a inhomogeneous medium (Inoue et al., 2011; Tomita et al., 2022), and lobes of radio galaxies (Matthews et al., 2019).
@@
To understand the dissipation in a collisionless shock in a relativistically hot plasma, as a first step, we investigated collisionless shocks propagating with the nonrelativistic and mildly relativistic velocity in a relativistically hot unmagnetized electron-positron plasma using two-dimensional particle-in-cell (PIC) simulations (Kamiido & Ohira, 2024). Our simulation showed that shocks are mediated by the Weibel instability, the strength of the Weibel magnetic fields becomes smaller as the shock velocity becomes smaller, and particle acceleration by the shock is not observed in the simulation time. In Kamiido & Ohira (2024), we discussed the reason for this shock velocity dependence in terms of the dissipation efficiency at the shock front, but did not discuss it in terms of the plasma instability. In addition, the temperature dependence and the case with a relativistic shock velocity were not investigated. In this study, we investigate collisionless shocks propagating with a relativistic shock velocity in relativistically hot unmagnetized electron-positron plasmas. The simulation setup and results are presented in Sections 2 and 3, respectively. To understand the simulation results, we calculate the pressure anisotropy in the shock transition region in Section 4. The discussion and summary are given in Section 5. In the appendices A, B, and C, we derive some physical quantities for the relativistic plasma.
@@ 2 Simulation setup
$n_{\mathrm{up}}$ is the upstream number density at the downstream rest frame (simulation frame). $\langle\gamma\rangle$ is the mean Lorentz factor of particles in the simulation frame and given by (Equation A2)
@@ 3.2 Magnetic field distribution
Figure 3 shows the distribution of the $z$-component of the magnetic field at $t\omega_{\mathrm{p}}=2000$, which is normalized by the upstream energy density $n_{\mathrm{up}}\langle\gamma\rangle mc^{2}$. The maximum value and typical wavelength of the magnetic field are $B_{z}/\sqrt{8\pi n_{\mathrm{up}}\langle\gamma\rangle mc^{2}}\sim 0.3$ and $\lambda\sim 8\,c/\omega_{\mathrm{p}}$ in common for all simulations, which are almost identical to the results in the case of cold upstream (Spitkovsky, 2008b). In addition, the decay rate of the downstream magnetic field is also common to all simulations. Since the Weibel-like magnetic field structure is observed in all simulations, our simulations show that the relativistic unmagnetized collisionless shock is mediated by the Weibel instability, even though the upstream temperature is relativistically hot.
@@ 3.3 Energy spectrum
Figure 4 shows the energy spectrum in the downstream region of $800500<x/(c/\omega_{\mathrm{p}})<$ 800 at $t\omega_{\mathrm{p}}=2000$. As one can see, the energy spectra have the power-law tails with the spectral index of about 2.4 for all simulations, that is, our simulations show that, as in the early studies for the relativistic Weibel mediated shocks in cold plasmas (Spitkovsky, 2008b; Sironi et al., 2013), particles are accelerated by the relativistic unmagnetized collisionless shock in the relativistically hot electron-positron plasma.
@@ 4 Pressure anisotropy
The plasma in the shock transition region consists mainly of two components: the upstream plasma moving from upstream to downstream and the plasma leaking from downstream to upstream. As a simple system, we consider counter streaming plasmas with the bulk velocities in the $x$ direction ±$\pm\beta c$, and the same density and temperature, $n$ and $T$. Even for a realistic asymmetric beam system, the pressure anisotropy does not change significantly (see appendix E for details).
The momentum distribution of the counter streaming plasmas is assumed to be the shifted Maxwell-Jüttner distribution function (Equation B2). Then, the pressure anisotropy of the counter streaming plasmas is given by (Equation B3),
@@
Figure 6: Distribution of counter streaming plasmas in the four velocity space. The four velocities $\mathbf{u}$ are normalized by the mean value of the momentum of particles in the beam moving in $+x$ direction, $\langle u_{x,+}\rangle$.
The white lines show the cone of the relativistic beaming effect, $u_{y}/u_{x}=\pm 1/\Gamma$, where $\Gamma=20$ is the bulk Lorentz factor corresponding to the counter streaming velocity.
Figure 6 shows the momentum distribution in the $u_{x}-u_{y}$ plane of the counter streaming plasmas with $\Gamma=20$ and $T/mc^{2}=0.01,5,500$, where $u_{x}$ and $u_{y}$ are the four velocity in the $x$ and $y$ directions, and normalized by $\langle u_{x,+}\rangle$. $\langle u_{x,+}\rangle$ is the mean value of the momentum of particles, $u_{x}$, in the beam moving in $+x$ direction in this system (equation A). As shown in the middle and bottom panels in Figure 6, the normalized momentum distribution does not depend on the temperature for the relativistic temperature. Since almost all particles have velocities very close to the speed of light in the each plasma rest frame for the relativistic temperature, in the counter streaming frame, their angular distribution is concentrated in the flow direction in the angle of $\Gamma^{-1}$. This is the so-called relativistic beaming effect. Then, the angular distribution of the counter streaming plasmas is almost independent of the temperature. This is the reason why the pressure anisotropy does not depend on the relativistic temperature.
@@ Appendix A Plasma with a relativistic temperature
where $e$ and $m$ are the charge and mass of each particle. The exact expression of the plasma frequency in the plasma rest frame is given by the dispersion relation in the Vlasov equation system (Bergman & Eliasson, 2001),
$\omega_{\mathrm{p}}^{2}=G^{30}_{10}\left.\quantity(\left.\frac{\theta^{2}}{4}\right|^{2}_{-1/2,1,3/2})\right/2K_{2}(\theta),$ (A4)
where $G$ is the Meijer’s G function. Our definition of the plasma frequency is easily to understand due to using the effective mass and there is almost no difference between our definition and the exact expression in the plasma rest frame as shown in figure 7.
Figure 7: The exact plasma frequency in (Bergman & Eliasson, 2001) (black solid line) and our definition (red dashed line).
The mean 4-velocity, $\langle u_{x}\rangle$ in the laboratory frame is also given by
[…]
$=\Gamma\beta\quantity[\frac{4}{\theta}+\frac{K_{1}(\theta)}{K_{2}(\theta)}]\,.$ (A5)
@@ Appendix C Rankine-Hugoniot relation
Appendix D Center of momentum system of two assymmetric beams
In appendix E, we consider the pressure anisotropy of two assymetric beams. Because the pressure is defined in the center of momentum system, the three velocities of two beams in the center of momentum system and their Lorentz factors are required.
First of all, we consider two beams ($i=1,2$) having three velocity in the laboratory frame $\beta_{i}$ in the $+x$ direction, velocity Lorentz factor $\Gamma_{i}$, the density in the beam rest frame $n^{\prime}_{i}$, and the temperature $T_{i}=mc^{2}/\theta_{i}$. In the laboratory frame, the center of momentum system moves with the three velocity $\beta_{\mathrm{G}}$ and the Lorentz factor $\Gamma_{\mathrm{G}}$. In the center of momentum system, two beams move with $\beta_{i\mathrm{G}}$ and $\Gamma_{i\mathrm{G}}$. Using the equation (A), in the center of momentum system, where the total momentum is zero, the equation holds as
$\Gamma_{1\mathrm{G}}^{2}\beta_{1\mathrm{G}}=r\Gamma_{2\mathrm{G}}^{2}\beta_{2\mathrm{G}}\,,$ (D1)
where
$r=\frac{n_{2}^{\prime}}{n_{1}^{\prime}}\frac{4/\theta_{2}+K_{1}(\theta_{2})/K_{2}(\theta_{2})}{4/\theta_{1}+K_{1}(\theta_{1})/K_{2}(\theta_{1})}\,.$ (D2)
The Lorentz transformations of the three velocity and the Lorentz factor are
$\beta_{i\mathrm{G}}$ $=\frac{\beta_{i}-\beta_{\mathrm{G}}}{1-\beta_{i}\beta_{\mathrm{G}}}\,,$ (D3)
$\Gamma_{i\mathrm{G}}$ $=\Gamma_{i}\Gamma_{\mathrm{G}}(1-\beta_{i}\beta_{\mathrm{G}})\,.$ (D4)
Using these equations, $\beta_{\mathrm{G}}$ and $\Gamma_{\mathrm{G}}$ are derived as
$\beta_{\mathrm{G}}$ $=\frac{b-\sqrt{b^{2}-4a^{2}}}{2a}\,,$ (D5)
$\Gamma_{\mathrm{G}}^{2}$ $=\frac{b+\sqrt{b^{2}-4a^{2}}}{2\sqrt{b^{2}-4a^{2}}}\,,$ (D6)
where
$a$ $=\Gamma_{1}^{2}\beta_{1}+r\Gamma_{2}^{2}\beta_{2}\,,$ (D7)
$b$ $=\Gamma_{1}^{2}(1+\beta^{2}_{1})+r\Gamma_{2}^{2}(1+\beta^{2}_{2})\,,$ (D8)
$b^{2}-4a^{2}$ $=(r+1)^{2}+4r\Gamma_{1}^{2}\Gamma_{2}^{2}(\beta_{1}-\beta_{2})^{2}\,.$ (D9)
Finally, the expression of quantities in the center of momentum system is derived as
$\beta_{i\mathrm{G}}$ $=\frac{2a\beta_{i}-(b-\sqrt{b^{2}-4a^{2}})}{2a-\beta_{i}(b-\sqrt{b^{2}-4a^{2}})}\,,$ (D10)
$\Gamma_{i\mathrm{G}}$ $=\Gamma_{i}\quantity(\frac{b+\sqrt{b^{2}-4a^{2}}}{2\sqrt{b^{2}-4a^{2}}})^{1/2}\quantity(1-\beta_{i}\frac{b-\sqrt{b^{2}-4a^{2}}}{2a})$ (D11)
These become the simple expression when the extreme cases. When $r\rightarrow 0$, $\Gamma_{1\mathrm{G}}=1$, and $\Gamma_{2\mathrm{G}}=\Gamma_{2}\Gamma_{1}(1-\beta_{2}\beta_{1})$. When $\beta_{2}=-\beta_{1}=-1$ and $r^{1/2},\,r^{-1/2}\ll\Gamma_{1}^{2}\Gamma_{2}^{2}$, $\Gamma_{1\mathrm{G}}=r^{1/4}\sqrt{\Gamma_{1}\Gamma_{2}}$ and $\Gamma_{2\mathrm{G}}=r^{-1/4}\sqrt{\Gamma_{1}\Gamma_{2}}$.
Appendix E Pressure anisotropy of two asymmetric beams
In section 4, we discuss the pressure anisotropy in the transition layer using the very simple symmetric system, i.e. counter streaming plasmas. After time development, the momentum distribution in the shock transition layer becomes asymmetric since there are two components, the dense upstream component and the leakage component from the downstream. In this section, we show such the asymmetric system is essentially the same as the symmetric system as shown in section 4, and our discussion using the symmetric system is suffice.
We consider two beams in the laboratory frame. The momentum of two beams having the number density in the beam rest frame $n^{\prime}_{i}$, the 3-velocity $\beta_{i}$, the Lorentz factor $\Gamma_{i}$, and the temperature $T_{i}=mc^{2}/\theta_{i}$. Because the pressure is defined in the center of momentum system, we have to move to the center of momentum system. As calculated in appendix B,
$P_{xx}$ $=\Gamma_{1\mathrm{G}}n_{1}^{\prime}\quantity[\frac{1}{\theta_{1}}(1+3\beta_{1\mathrm{G}}^{2})+\beta_{1\mathrm{G}}^{2}\frac{K_{1}(\theta_{1})}{K_{2}(\theta_{1})}]$
$~{}~{}~{}+\Gamma_{2\mathrm{G}}n_{2}^{\prime}\quantity[\frac{1}{\theta_{2}}(1+3\beta_{2\mathrm{G}}^{2})+\beta_{2\mathrm{G}}^{2}\frac{K_{1}(\theta_{2})}{K_{2}(\theta_{2})}]\,,$ (E1)
and
$P_{yy}=\frac{n_{1}^{\prime}}{\theta_{1}}+\frac{n_{2}^{\prime}}{\theta_{2}}\,.$ (E2)
$\beta_{i\mathrm{G}}$ and $\Gamma_{i\mathrm{G}}$ is calculated using equations (D10), (D11). When we assume $n_{1}^{\prime}=n_{2}^{\prime}$, $\beta_{1}=-\beta_{2}$, and $\theta_{1}=\theta_{2}$, we can get the same pressure anisotropy in appendix B.
In the shock transition layer, two components exists (index $\mathrm{u}$ : upstream fluid and index $\mathrm{l}$ : leakage from the downstream). Considering the extreme case, $\beta_{\mathrm{l}}\sim 1$, $\beta_{\mathrm{u}}\sim-1$, $\Gamma_{\mathrm{u}}=\Gamma_{\mathrm{l}}=\Gamma$, $\theta_{\mathrm{u}}=\theta_{\mathrm{l}}=\theta$, and $n_{\mathrm{u}}^{\prime}/n_{\mathrm{l}}^{\prime},\,n_{\mathrm{l}}^{\prime}/n_{\mathrm{u}}^{\prime}\ll\Gamma^{8}$, the pressure anisotropy is calculated as
$\frac{P_{xx}}{P_{yy}}-1=\Gamma^{2}\frac{2\sqrt{n_{\mathrm{l}}^{\prime}/n_{\mathrm{u}}^{\prime}}}{1+n_{\mathrm{l}}^{\prime}/n_{\mathrm{u}}^{\prime}}\quantity[4+\theta\frac{K_{1}(\theta)}{K_{2}(\theta)}]-1\,.$ (E3)
As long as $n_{\mathrm{l}}^{\prime}/n_{\mathrm{u}}^{\prime}\sim\order{0.1}$ ), the above estimation does not significantly differ from one for the symmetric beam system as shown in section 4.